OpenStudy (anonymous):

Differentiate 2x^2-6x^2+y^3=26 with respect to t (x and y are functions of t)

OpenStudy (anonymous):

take the derivative and put \[x'=\frac{dx}{dt}\] or \[y'=\frac{dy}{dt}\] after each one.

OpenStudy (anonymous):

but first, so you really have \[2x^2-6x^2=-4x^2\]as the first two terms?

OpenStudy (anonymous):

4x(dx/dt)-12x(dx/dt)+3y^2(dy/dt)=0

OpenStudy (anonymous):

so I solve for y? and then find the derivative?

OpenStudy (anonymous):

\[4x * x \prime - 12y* y \prime + 3y2*y \prime?\]